OpenStudy (anonymous):

snookers lumbers can convert logs into either lumber or plywood in a given day the mill turns out twice as many units of plywood as lumber it makes a profit of $20 on a unit of lumber and $40 on a unit of plywood how many of each unit must be produced and sold in order to make a profit of $10100

OpenStudy (anonymous):

Let the number of units of lumber = x Then number of units of plywood = 2x Profit earned on x units of lumber = 20 * x = 20x Profit earned on 2x units of plywood = 40 * 2x = 80x ATQ, 20x + 80x = 10100 100x = 10100 x = 10100/100 = 101 Therefore, number of units of lumber is 101 and Number of units of plywood is 202 CHECKING : 101 x 20 = 2020 202 x 40 = 8080 ------ 10100 Hope this clear. If yes, pls click GOOD ANSWER button to close the question

OpenStudy (anonymous):

thanks harkirat can you help with this solve by elimination method 5x+2y=-13 7x-3y=19

OpenStudy (anonymous):

Hi Nancy can you help with this one solve by elimination method 5x+2y=-13 7x-3y=19

OpenStudy (anonymous):

Multiply first equation by 7 and second by 5 to get 35x + 14y = -91 ------- (A) 35x - 15y = 95 ------- (B)

OpenStudy (anonymous):

Now subtract (B) frm (A) to get 29y = - 186 so y = - 186/29 Putting y = -186/29 in the given first equation 5x + 2y = -13, we get 5x + 2 (-186/29) = -13 5x - 372/29 = -13 5x = -13 + 372/29 5x = (-377 + 372)/29 5x = -5/29 x = -5/ (29 * 5) x = -1/29

OpenStudy (anonymous):

jazil, is it clear???

OpenStudy (anonymous):

yes thank you very much

OpenStudy (anonymous):

click good answer pls